# Croppable Knowledge Graph Embedding

### v1

@@ 5.6.1 Training efficiency
TransE RotatE PairRE
WN DT 74.0 (9.49$\times$) 141.0 (11.10$\times$) 67.4 (10.06$\times$)
Ext-based 1.5 (0.19$\times$) 2.5 (0.20$\times$) 1.6 (0.24$\times$)
BKD 91.5 (11.73$\times$) 163.0 (12.83$\times$) 87.5 (13.06$\times$)
TA 172.0 (22.05$\times$) 272.0 (21.42$\times$) 166.0 (24.78$\times$)
DualDE 151.0 (19.36$\times$) 240.0 (18.90$\times$) 133.0 (19.85$\times$)
MED 7.8 (1.00$\times$) 12.7 (1.00$\times$) 6.7 (1.00$\times$)
FB DT 218.0 (10.23$\times$) 381.0 (10.73$\times$) 179.0 (9.37$\times$)
Ext-based 4.7 (0.22$\times$) 9.5 (0.27$\times$) 3.7 (0.19$\times$)
BKD 248.0 (11.64$\times$) 443.0 (12.48$\times$) 231.0 (12.09$\times$)
TA - - - - - -
DualDE - - - - - -
MED 21.3 (1.00$\times$) 35.5 (1.00$\times$) 19.1 (1.00$\times$)
We report the training time of obtaining 64 models of all sizes ($d$=10, 20, …, 640) by different methods in Table 7. For DT, the training time cost is the sum of the time of directly training 64 KGE models of all sizes in turn. For the Ext-based baselines, the training time cost is the same and is equal to the time of training a $d_{n}$-dimensional KGE model since the time of arranging dimensions is very short and negligible. For the KD-based baselines, the training time cost is the sum of the time of training the $d_{n}$-dimensional teacher model and distilling 63 student models ($d$=10, 20, …, 630) in turn. All training is performed on a single NVIDIA Tesla A100 40GB GPU for fair comparison. For TA and DualDE on FB15K237, we don’t train student models of all 63 sizes, which is estimated to take more than 400 hours on each KGE method. Compared with directly trained (DT) models of all sizes in turn, MED accelerates by up to 10$\times$ for 3 KGE methods. Although Ext-based baselines spend the shortest training time, they perform particularly poorly and lack practical value. TA and DualDE need to optimize both the student model and a larger teacher model, which greatly increases the training parameters and time cost.
5.6.2 Whether high-dimensional sub-models cover the capabilities of low-dimensional ones
If a high-dimensional model retains the ability of lower-dimensional models, it should correctly predict all triples that the lower-dimensional model can predict. We count the percentage of triples in test set that meet the condition that if the smallest sub-model that can correctly predict a given triple is $M_{i}$, all higher-dimensional sub-models ($M_{i+1}$, $M_{i+2}$, …, $M_{n}$) also correctly predict it, and denote the result as the ability retention ratio (ARR). We use Hit@10 to judge whether a triple is correctly predicted, that is, $M_{i}$ correctly predicts a triple if $M_{i}$ scores this triple in the top 10 among all candidate triples.
@@ 5.6.2 Whether high-dimensional sub-models cover the capabilities of low-dimensional ones
From Fig. 4, we find that ARR of MED is always much higher than baselines, especially on FB15K237, indicating that high-dimensional sub-models in MED successfully cover the power of low-dimensional ones, contributed by the mutual learning mechanism that helps high-dimensional sub-models review what low-dimensional sub-models have learned. Based on this advantage of MED, we can also provide a simple way to judge how easy or difficult a triple is for KGE methods to learn: the triple that low-dimensional sub-models can correctly predict may be easy since more high-dimensional models can also predict it, while triples that can only be predicted by a particularly high-dimensional sub-model are difficult.
5.6.3 Visual analysis of embedding
Figure 5: Clustering on FB15K237 with RotatE.
We select four primary entity categories (‘organization’, ‘sports’, ‘location’, and ‘music’) that contain more than 300 entities in FB15K237, and randomly select 250 entities for each. We cluster these entities’ embeddings of 3 different dimensions ($d$=10, 100, 600) by the t-SNE algorithm, and the clustering results are visualized in Fig. 5. Under the same dimension, the clustering result of MED is always the best, followed by DualDE, while the result of Ext-V is generally poor, which is consistent with the conclusion in Section 5.2. We also find some special phenomenons for MED when dimension increases: 1) the nodes of the ‘sports’ gradually become two clusters meaning MED learns more fine-grained category information as dimension increases. and 2) the relative distribution among different categories hardly changes and shows a trend of “inheritance” and “improvement”. This further proves MED achieves our expectation that high-dimensional sub-models retain the ability of low-dimensional sub-models, and can learn more knowledge than low-dimensional sub-models.
@@ 6 Conclusion
In this work, we propose a novel KGE training framework, MED, that trains a croppable KGE at once, and then sub-models of various required dimensions can be cropped out from it and used directly without additional training. In MED, we propose the mutual learning mechanism to improve low-dimensional sub-models performance and make the high-dimensional sub-models retain the ability of the low-dimensional ones, the evolutionary improvement mechanism to motivate high-dimensional sub-models to master more knowledge that low-dimensional ones cannot, and the dynamic loss weight to adaptively balance multiple losses. The experimental results show the effectiveness and high efficiency of our method, where all sub-models achieve promising performance, especially the performance of low-dimensional sub-models is greatly improved. In future work, we will further explore the more fine-grained information encoding ability of each sub-model.
@@ Appendix A More Results of link prediction
Method MRR Hit@1 MRR Hit@1 MRR Hit@1 MRR Hit@1 MRR Hit@1 MRR Hit@1 MRR Hit@1
TransE DT .121 .011 .176 .016 .214 .018 .227 .025 .233 .027 .235 .033 .237 .034
Ext .125 .016 .172 .023 .199 .023 .213 .028 .225 .033 .226 .028 .237 .034
Ext-L .139 .029 .196 .025 .224 .039 .232 .046 .236 .036 .236 .033 .237 .034
Ext-V .139 .029 .198 .045 .222 .051 .234 .047 .236 .036 .236 .027 .237 .034
BKD .141 .035 .207 .040 .226 .033 .232 .031 .233 .030 .236 .032 - -
TA .144 .040 .211 .043 .226 .037 .233 .030 .234 .030 .236 .034 - -
DualDE .148 .037 .213 .043 .225 .037 .234 .031 .235 .031 .238 .034 - -
MED .170 .040 .219 .045 .232 .048 .232 .042 .236 .037 .237 .033 .237 .031
RotatE DT .172 .005 .409 .357 .456 .393 .465 .420 .471 .423 .474 .428 .476 .429
Ext .299 .257 .379 .335 .437 .395 .458 .415 .467 .413 .471 .418 .476 .429
Ext-L .206 .166 .336 .288 .399 .352 .423 .373 .445 .396 .466 .417 .476 .429
Ext-V .261 .197 .304 .234 .337 .263 .366 .293 .416 .357 .451 .397 .476 .429
BKD .175 .009 .424 .361 .457 .403 .471 .421 .472 .424 .474 .425 - -
TA .177 .010 .424 .363 .459 .408 .470 .420 .473 .422 .474 .425 - -
DualDE .179 .011 .425 .364 .462 .412 .471 .423 .473 .426 .475 .425 - -
MED .324 .277 .456 .409 .466 .418 .471 .422 .471 .424 .476 .427 .476 .428
PairRE DT .220 .174 .342 .313 .415 .384 .435 .399 .449 .405 .452 .406 .453 .407
Ext .152 .120 .261 .198 .334 .267 .375 .314 .419 .364 .438 .388 .453 .407
Ext-L .162 .129 .281 .237 .363 .319 .417 .377 .437 .395 .446 .400 .453 .407
Ext-V .172 .124 .306 .269 .389 .352 .420 .379 .441 .398 .446 .400 .453 .407
BKD .228 .184 .375 .334 .421 .372 .443 .405 .451 .405 .453 .407 - -
TA .245 .197 .381 .332 .426 .380 .448 .404 .452 .409 .453 .408 - -
DualDE .242 .175 .377 .330 .428 .381 .451 .409 .453 .410 .454 .410 - -
MED .317 .259 .408 .367 .433 .392 .449 .405 .451 .406 .451 .407 .451 .406
Table 9: Hit@10 and Hit@3 of some representative dimensions on WN18RR.
@@
Method Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3
TransE DT .287 .202 .453 .291 .496 .385 .524 .401 .531 .403 .534 .407 .537 .412
Ext .298 .201 .423 .285 .468 .338 .495 .364 .515 .384 .521 .388 .537 .412
Ext-L .315 .218 .461 .317 .497 .361 .516 .403 .534 .405 .535 .408 .537 .412
Ext-V .309 .218 .458 .314 .494 .391 .525 .407 .532 .408 .536 .411 .537 .412
BKD .323 .216 .480 .331 .513 .392 .527 .401 .531 .404 .533 .407 - -
TA .335 .224 .483 .343 .512 .395 .527 .408 .533 .407 .535 .410 - -
DualDE .337 .226 .488 .346 .514 .394 .530 .408 .533 .408 .535 .411 - -
MED .388 .269 .491 .369 .518 .399 .523 .404 .529 .407 .536 .410 .537 .412
RotatE DT .418 .304 .504 .436 .556 .475 .564 .487 .567 .489 .573 .491 .575 .493
Ext .378 .315 .464 .399 .516 .452 .544 .472 .549 .480 .552 .470 .575 .493
Ext-L .277 .224 .424 .359 .487 .420 .515 .441 .541 .461 .564 .481 .575 .493
Ext-V .377 .289 .433 .336 .471 .377 .497 .402 .532 .442 .561 .467 .575 .493
BKD .434 .312 .540 .452 .556 .479 .565 .487 .570 .490 .572 .492 - -
TA .438 .314 .542 .452 .558 .481 .567 .489 .572 .488 .572 .492 - -
DualDE .440 .320 .542 .452 .559 .483 .567 .489 .573 .488 .573 .491 - -
MED .469 .354 .543 .476 .561 .486 .568 .490 .574 .492 .573 .493 .574 .495
PairRE DT .321 .271 .381 .368 .472 .428 .516 .450 .534 .463 .542 .462 .544 .464
Ext .209 .163 .379 .292 .463 .366 .493 .398 .526 .437 .545 .452 .544 .464
Ext-L .220 .175 .360 .302 .442 .383 .495 .431 .523 .450 .544 .455 .544 .464
Ext-V .260 .192 .374 .323 .456 .407 .498 .435 .529 .452 .541 .458 .544 .464
BKD .336 .279 .413 .388 .483 .435 .525 .452 .536 .460 .542 .463 - -
TA .340 .293 .427 .387 .487 .437 .534 .460 .537 .462 .543 .463 - -
DualDE .336 .281 .424 .389 .495 .437 .536 .463 .540 .463 .544 .465 - -
MED .376 .314 .467 .426 .502 .443 .537 .462 .541 .464 .542 .465 .542 .464
Table 10: MRR and Hit@1 of some representative dimensions on FB15K237.

### v2

@@ 5.6.1 Training efficiency
TransE SimplE RotatE PairRE
WN DT 74.0 (9.49$\times$) 68.0 (12.14$\times$) 141.0 (11.10$\times$) 67.4 (10.06$\times$)
Ext-based 1.5 (0.19$\times$) 1.3 (0.23$\times$) 2.5 (0.20$\times$) 1.6 (0.24$\times$)
BKD 91.5 (11.73$\times$) 72.0 (12.86$\times$) 163.0 (12.83$\times$) 87.5 (13.06$\times$)
TA 172.0 (22.05$\times$) 142.0 (25.36$\times$) 272.0 (21.42$\times$) 166.0 (24.78$\times$)
DualDE 151.0 (19.36$\times$) 133.0 (23.75$\times$) 240.0 (18.90$\times$) 133.0 (19.85$\times$)
IterDE 140.9 (18.06$\times$) 118.0 (21.07$\times$) 216.0 (17.01$\times$) 124.0 (18.51$\times$)
MED 7.8 (1.00$\times$) 5.6 (1.00$\times$) 12.7 (1.00$\times$) 6.7 (1.00$\times$)
FB DT 218.0 (10.23$\times$) 179.0 (10.65$\times$) 381.0 (10.73$\times$) 179.0 (9.37$\times$)
Ext-based 4.7 (0.22$\times$) 5.1 (0.30$\times$) 9.5 (0.27$\times$) 3.7 (0.19$\times$)
BKD 248.0 (11.64$\times$) 227.0 (13.51$\times$) 443.0 (12.48$\times$) 231.0 (12.09$\times$)
MED 21.3 (1.00$\times$) 16.8 (1.00$\times$) 35.5 (1.00$\times$) 19.1 (1.00$\times$)
@@
Figure 4: Sub-models’ MRR during training on WN18RR with RotatE.
Table 7 reports the training time of getting 64 models of various sizes ($d$=10, 20, …, 640). Figure 4 shows MRR change of MED’s sub-models during training. For fair comparison, all training is done on a single NVIDIA Tesla A100 40GB GPU. For DT, the training time is the sum of directly training 64 KGE models sequentially. Ext-based baselines have the same training time, equal to that of training a $d_{n}$-dimensional KGE model as dimension arrangement time is negligible. KD-based baselines’ training time is the sum of training a $d_{n}$-dimensional teacher model and distilling 63 student models. On FB15K237, we don’t train all 63 sizes of student models for TA, DualDE, and IterDE, estimated to take over 400 hours per KGE method. Compared to DT, MED speeds up by up to 10$\times$ for 4 KGE methods. Ext-based baselines have the shortest training time but poor performance and low practical value. Except for BKD, other KD-based methods need to optimize both student and teacher, increasing train cost.
5.6.2 Effect of the number of sub-models
10d 40d 160d 640d
n train time MRR H10 MRR H10 MRR H10 MRR H10
64 12.7h 0.324 0.469 0.466 0.561 0.471 0.574 0.476 0.574
16 6.2h 0.322 0.467 0.465 0.561 0.473 0.575 0.477 0.576
4 3.3h 0.319 0.463 0.463 0.561 0.475 0.577 0.480 0.578
Table 8: Results of different $n$.
We set the number of different sub-models ($n$= 64, 16, 4) for RotatE on WN18RR. Table 8 shows that reducing the number of sub - models boosts high-dimensional ($d$=160 and 640) model performance but lowers that of low-dimensional ($d$=10 and 40) models.Training efficiency is nearly linearly related to the number of models.
5.6.3 Whether high-dimensional sub-models cover low-dimensional ones’ capabilities
A high-dimensional model retaining lower-dimensional ones’ ability should correctly predict all triples the latter can. We calculate the percentage of test-set triples where if the smallest sub-model correctly predicting a triple is $M_{i}$, all higher-dimensional ones ($M_{i+1}$, $M_{i+2}$, …, $M_{n}$) do too. This result is the ability retention ratio (ARR). We use Hit@10 to judge correct prediction: $M_{i}$ predicts a triple correctly if it scores in the top 10 among candidates.
@@ 5.6.3 Whether high-dimensional sub-models cover low-dimensional ones’ capabilities
Fig. 5 shows that MED’s ARR is is much higher than baselines, especially on FB15K237. This means MED’s high-dimensional sub-models cover low-dimensional ones’ power, thanks to the mutual learning mechanism for knowledge review. This advantage of MED provides a simple way to judge triple difficulty for KGE methods: triples low-dimensional sub-models predict may be easy as high-dimensional ones can too, while those only predicted by high-dimensional sub-models are difficult.
@@ 6 Conclusion
In this work, we propose a novel KGE training framework, MED, which trains a croppable KGE in one go. Sub-models of various dimensions can be cropped from it and used directly without extra training. In MED, we introduce the mutual learning mechanism to improve low-dimensional sub-models and make the high-dimensional sub-models retain low-dimensional ones’ ability, the evolutionary improvement mechanism to prompt high-dimensional sub-models to learn what low-dimensional ones can’t, and dynamic loss weights to balance multiple losses adaptively. Experimental results demonstrate MED’s effectiveness, high efficiency, and flexible extensibility.
7 Limitations
As a research paper, we are unable to conduct experiments on all knowledge graph embedding (KGE) methods. To demonstrate the generality of the method proposed in this paper, we selected representative models from different types of KGE methods, including distance-based TransE, rotation-based RotatE, semantic matching-based SimplE, and subrelation encoding-based PairRE. In fact, there are many KGE methods that outperform those used in our experiments. However, the method proposed in this paper is applicable to any KGE method with a triple scoring function.
8 Acknowledgements
This work is founded by National Natural Science Foundation of China (NSFCU23B2055/NSFCU19B2027/NSFC62306276), Zhejiang Provincial Natural Science Foundation of China (No. LQ23F020017), Yongjiang Talent Introduction Programme (2022A-238-G), and Fundamental Research Funds for the Central Universities (226-2023-00138). This work was supported by AntGroup.
@@ Appendix A More Results of link prediction
TransE Method MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
DT 0.121 0.287 0.176 0.453 0.214 0.496 0.227 0.524 0.233 0.531 0.235 0.534 0.237 0.537
Ext 0.125 0.298 0.172 0.423 0.199 0.468 0.213 0.495 0.225 0.515 0.226 0.521 0.237 0.537
Ext-L 0.139 0.315 0.196 0.461 0.224 0.497 0.232 0.516 0.236 0.534 0.236 0.535 0.237 0.537
Ext-V 0.139 0.309 0.198 0.458 0.222 0.494 0.234 0.525 0.236 0.532 0.236 0.536 0.237 0.537
BKD 0.141 0.323 0.207 0.480 0.226 0.513 0.232 0.527 0.233 0.531 0.236 0.533 - -
TA 0.144 0.335 0.211 0.483 0.226 0.512 0.233 0.527 0.234 0.533 0.236 0.535 - -
DualDE 0.148 0.337 0.213 0.488 0.225 0.514 0.234 0.530 0.235 0.533 0.238 0.535 - -
IterDE 0.143 0.332 0.211 0.484 0.224 0.511 0.232 0.528 0.236 0.531 0.237 0.533 - -
MED 0.170 0.388 0.219 0.491 0.232 0.518 0.232 0.523 0.236 0.529 0.237 0.536 0.237 0.537
SimplE Method MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
DT 0.061 0.126 0.257 0.372 0.316 0.389 0.382 0.446 0.409 0.459 0.417 0.474 0.421 0.481
Ext 0.004 0.007 0.051 0.107 0.160 0.249 0.219 0.314 0.357 0.401 0.407 0.451 0.421 0.481
Ext-L 0.005 0.006 0.048 0.078 0.169 0.244 0.369 0.435 0.398 0.454 0.417 0.481 0.421 0.481
Ext-V 0.004 0.006 0.047 0.076 0.246 0.317 0.368 0.402 0.398 0.461 0.413 0.472 0.421 0.481
BKD 0.075 0.156 0.285 0.381 0.343 0.399 0.394 0.450 0.414 0.468 0.418 0.475 - -
TA 0.089 0.189 0.316 0.386 0.368 0.418 0.405 0.456 0.415 0.472 0.421 0.481 - -
DualDE 0.083 0.175 0.328 0.388 0.386 0.423 0.407 0.454 0.419 0.475 0.422 0.482 - -
IterDE 0.077 0.162 0.321 0.378 0.375 0.419 0.404 0.452 0.416 0.469 0.421 0.482 - -
MED 0.111 0.224 0.335 0.395 0.385 0.431 0.407 0.457 0.418 0.477 0.421 0.481 0.421 0.482
RotatE Method MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
DT 0.172 0.418 0.409 0.504 0.456 0.556 0.465 0.564 0.471 0.567 0.474 0.573 0.476 0.575
Ext 0.299 0.378 0.379 0.464 0.437 0.516 0.458 0.544 0.467 0.549 0.471 0.552 0.476 0.575
Ext-L 0.206 0.277 0.336 0.424 0.399 0.487 0.423 0.515 0.445 0.541 0.466 0.564 0.476 0.575
Ext-V 0.261 0.377 0.304 0.433 0.337 0.471 0.366 0.497 0.416 0.532 0.451 0.561 0.476 0.575
BKD 0.175 0.434 0.424 0.540 0.457 0.556 0.471 0.565 0.472 0.570 0.474 0.572 - -
TA 0.177 0.438 0.424 0.542 0.459 0.558 0.470 0.567 0.473 0.572 0.474 0.572 - -
DualDE 0.179 0.440 0.425 0.542 0.462 0.559 0.471 0.567 0.473 0.573 0.475 0.573 - -
IterDE 0.176 0.436 0.421 0.538 0.459 0.560 0.470 0.567 0.471 0.569 0.474 0.572 - -
MED 0.324 0.469 0.456 0.543 0.466 0.561 0.471 0.568 0.471 0.574 0.476 0.573 0.476 0.574
PairRE Method MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
DT 0.220 0.321 0.342 0.381 0.415 0.472 0.435 0.516 0.449 0.534 0.452 0.542 0.453 0.544
Ext 0.152 0.209 0.261 0.379 0.334 0.463 0.375 0.493 0.419 0.526 0.438 0.545 0.453 0.544
Ext-L 0.162 0.220 0.281 0.360 0.363 0.442 0.417 0.495 0.437 0.523 0.446 0.544 0.453 0.544
Ext-V 0.172 0.260 0.306 0.374 0.389 0.456 0.420 0.498 0.441 0.529 0.446 0.541 0.453 0.544
BKD 0.228 0.336 0.375 0.413 0.421 0.483 0.443 0.525 0.451 0.536 0.453 0.542 - -
TA 0.245 0.340 0.381 0.427 0.426 0.487 0.448 0.534 0.452 0.537 0.453 0.543 - -
DualDE 0.242 0.336 0.377 0.424 0.428 0.495 0.451 0.536 0.453 0.540 0.454 0.544 - -
IterDE 0.235 0.336 0.379 0.423 0.426 0.495 0.449 0.533 0.450 0.538 0.452 0.543 - -
MED 0.317 0.376 0.408 0.467 0.433 0.502 0.449 0.537 0.451 0.541 0.451 0.542 0.451 0.542
Table 9: MRR and Hit@10 of some representative dimensions on WN18RR.
@@
TransE Method Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1
DT 0.202 0.011 0.291 0.016 0.385 0.018 0.401 0.025 0.403 0.027 0.407 0.033 0.412 0.034
Ext 0.201 0.016 0.285 0.023 0.338 0.023 0.364 0.028 0.384 0.033 0.388 0.028 0.412 0.034
Ext-L 0.218 0.029 0.317 0.025 0.361 0.039 0.403 0.046 0.405 0.036 0.408 0.033 0.412 0.034
Ext-V 0.218 0.029 0.314 0.045 0.391 0.051 0.407 0.047 0.408 0.036 0.411 0.027 0.412 0.034
BKD 0.216 0.035 0.331 0.040 0.392 0.033 0.401 0.031 0.404 0.030 0.407 0.032 - -
TA 0.224 0.040 0.343 0.043 0.395 0.037 0.408 0.030 0.407 0.030 0.410 0.034 - -
DualDE 0.226 0.037 0.346 0.043 0.394 0.037 0.408 0.031 0.408 0.031 0.411 0.034 - -
IterDE 0.217 0.032 0.345 0.044 0.392 0.036 0.407 0.030 0.408 0.031 0.407 0.033 - -
MED 0.269 0.040 0.369 0.045 0.399 0.038 0.404 0.042 0.407 0.037 0.410 0.033 0.412 0.031
SimplE Method Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1
DT 0.061 0.028 0.297 0.193 0.361 0.289 0.406 0.343 0.420 0.382 0.428 0.386 0.433 0.391
Ext 0.003 0.001 0.055 0.023 0.181 0.114 0.249 0.168 0.377 0.329 0.422 0.381 0.433 0.391
Ext-L 0.004 0.003 0.051 0.031 0.187 0.128 0.389 0.333 0.413 0.365 0.429 0.384 0.433 0.391
Ext-V 0.004 0.002 0.050 0.029 0.269 0.205 0.378 0.349 0.409 0.372 0.426 0.382 0.433 0.391
BKD 0.077 0.034 0.331 0.225 0.384 0.311 0.415 0.358 0.426 0.371 0.431 0.385 - -
TA 0.093 0.042 0.349 0.269 0.375 0.349 0.412 0.384 0.425 0.388 0.431 0.389 - -
DualDE 0.086 0.038 0.361 0.285 0.391 0.368 0.416 0.383 0.427 0.389 0.434 0.392 - -
IterDE 0.079 0.033 0.355 0.279 0.382 0.356 0.415 0.379 0.424 0.383 0.433 0.389 - -
MED 0.119 0.048 0.366 0.292 0.395 0.359 0.419 0.380 0.429 0.389 0.435 0.391 0.434 0.390
RotatE Method Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1
DT 0.304 0.005 0.436 0.357 0.475 0.393 0.487 0.420 0.489 0.423 0.491 0.428 0.493 0.429
Ext 0.315 0.257 0.399 0.335 0.452 0.395 0.472 0.415 0.480 0.413 0.470 0.418 0.493 0.429
Ext-L 0.224 0.166 0.359 0.288 0.420 0.352 0.441 0.373 0.461 0.396 0.481 0.417 0.493 0.429
Ext-V 0.289 0.197 0.336 0.234 0.377 0.263 0.402 0.293 0.442 0.357 0.467 0.397 0.493 0.429
BKD 0.312 0.009 0.452 0.361 0.479 0.403 0.487 0.421 0.490 0.424 0.492 0.425 - -
TA 0.314 0.010 0.452 0.363 0.481 0.408 0.489 0.420 0.488 0.422 0.492 0.425 - -
DualDE 0.320 0.011 0.452 0.364 0.483 0.412 0.489 0.423 0.488 0.426 0.491 0.425 - -
IterDE 0.311 0.013 0.439 0.356 0.479 0.407 0.484 0.423 0.488 0.425 0.493 0.424 - -
MED 0.354 0.277 0.476 0.409 0.486 0.418 0.490 0.422 0.492 0.424 0.493 0.427 0.495 0.428
PairRE Method Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1
DT 0.271 0.174 0.368 0.313 0.428 0.384 0.450 0.399 0.463 0.405 0.462 0.406 0.464 0.407
Ext 0.163 0.120 0.292 0.198 0.366 0.267 0.398 0.314 0.437 0.364 0.452 0.388 0.464 0.407
Ext-L 0.175 0.129 0.302 0.237 0.383 0.319 0.431 0.377 0.450 0.395 0.455 0.400 0.464 0.407
Ext-V 0.192 0.124 0.323 0.269 0.407 0.352 0.435 0.379 0.452 0.398 0.458 0.400 0.464 0.407
BKD 0.279 0.184 0.388 0.334 0.435 0.372 0.452 0.405 0.460 0.405 0.463 0.407 - -
TA 0.293 0.197 0.387 0.332 0.437 0.380 0.460 0.404 0.462 0.409 0.463 0.408 - -
DualDE 0.281 0.175 0.389 0.330 0.437 0.381 0.463 0.409 0.463 0.410 0.465 0.410 - -
IterDE 0.285 0.172 0.390 0.331 0.435 0.377 0.461 0.405 0.463 0.411 0.464 0.410 - -
MED 0.314 0.259 0.426 0.367 0.443 0.392 0.462 0.405 0.464 0.406 0.465 0.407 0.464 0.406
Table 10: Hit@3 and Hit@1 of some representative dimensions on WN18RR.